### v1

# Memory-Enhanced Neural Solvers for Efficient Adaptation in Combinatorial Optimization

Abstract
Combinatorial Optimization is crucial to numerous real-world applications, yet still presents challenges due to its (NP-)hard nature. Amongst existing approaches, heuristics often offer the best trade-off between quality and scalability, making them suitable for industrial use. While Reinforcement Learning (RL) offers a flexible framework for designing heuristics, its adoption over handcrafted heuristics remains incomplete within industrial solvers. Existing learned methods still lack the ability to adapt to specific instances and fully leverage the available computational budget. The current best methods either rely on a collection of pre-trained policies, or on data-inefficient fine-tuning; hence failing to fully utilize newly available information within the constraints of the budget. In response, we present MEMENTO, an RL approach that leverages memory to improve the adaptation of neural solvers at inference time. MEMENTO enables updating the action distribution dynamically based on the outcome of previous decisions. We validate its effectiveness on benchmark problems, in particular Traveling Salesman and Capacitated Vehicle Routing, demonstrating it can successfully be combined with standard methods to boost their performance under a given budget, both in and out-of-distribution, improving their performance on all $12$ evaluated tasks.
****footnotetext: Work completed during an internship at InstaDeep\$$\dagger$\$$\dagger$\$footnotetext: Equal supervision
1 Introduction
Combinatorial Optimization (CO) encompasses a vast range of real-world applications, ranging from transportation (J-F Audy and Rousseau, 2011) and logistics (Dincbas et al., 1992) to energy management (Froger et al., 2016). These problems involve finding optimal orderings, labels, or subsets of discrete sets to optimize given objective functions. Real-world CO problems are typically NP-hard with a solution space growing exponentially with the problem size, making it intractable to find the optimal solution. Hence, industrial solvers rely on sophisticated heuristic approaches to solve them in practice. Reinforcement Learning (RL) provides a versatile framework for learning such heuristics and has demonstrated remarkable success in tackling CO tasks (Mazyavkina et al., 2021).
Traditionally, RL methods train policies to incrementally construct solutions. However, achieving optimality in a single construction attempt for NP-hard problems is impractical. Therefore, pre-trained policies are often combined with search procedures. The literature introduces a range of these procedures from stochastic sampling (Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023), beam search (Steinbiss et al., 1994), or Monte Carlo Tree Search (MCTS) (Browne et al., 2012) to more advanced strategies like online fine-tuning (Bello et al., 2016; Hottung et al., 2022) or searching a space of diverse pre-trained policies (Chalumeau et al., 2023b). One popular online fine-tuning strategy, Efficient Active Search (EAS) (Hottung et al., 2022), uses transitions from generated solutions to derive policy gradient updates. However, it suffers from the inherent drawbacks from back-propagation, in particular having each data point only impacting the update as much as what is enabled by the learning rate. The leading strategy, COMPASS (Chalumeau et al., 2023b), relies on a space of diverse pre-trained policies for fast adaptation, but is limited to selecting the most appropriate one, and lacks an update mechanism using the collected data. Meanwhile, the use of memory has grown in modern deep learning, demonstrated by the success of retrieval augmented approaches in Natural Language Processing (Lewis et al., 2020), and to a smaller extent, its use in RL (Humphreys et al., 2022). These mechanisms create a closer link between collected experience and policy update, making them a promising candidate to improve adaptation.
In this vein, we introduce MEMENTO, a method to efficiently update the action distribution of neural solvers online, leveraging a memory of recently collected data. Importantly, MEMENTO is model agnostic and can be applied to many existing neural solvers. MEMENTO is able to meta-learn expressive update rules, which prove experimentally to outperform classic policy gradient updates. We evaluate our method on two popular routing problems - Traveling Salesman (TSP) and Capacitated Vehicle Routing (CVRP). We evaluate all methods both in and out of distribution, and tackle instances up to size 500 to better understand their scaling properties. We experimentally show that MEMENTO can be combined with existing methods and provide systematic improvements, reaching state-of-the-art on $11$ out of $12$ tasks. We provide an analysis of the update rules learned by MEMENTO, which enables to adapt faster than classic policy gradient methods like EAS.
Our contributions come as follows: (i) We introduce MEMENTO, a flexible and scalable method – composed of a memory and a processing module – enabling efficient adaptation of policies at inference time. (ii) We provide experimental evidence that MEMENTO can be combined with existing approaches to boost their performance in and out of distribution, even for large instances. (iii) Whilst doing so, we train and evaluate leading construction methods on TSP and CVRP instances of size $500$ solely with RL, proving to outperform all existing RL methods, and release the checkpoints. (iv) We open-source the implementation of MEMENTO in JAX (Bradbury et al., 2018), along with test sets and checkpoints to facilitate comparison and future research advancements.
2 Related work
Construction methods for CO
Construction approaches in RL for CO refer to methods that incrementally build a solution one action after the other. Hopfield and Tank (1985) was the first to use a neural network to solve the TSP, followed by Bello et al. (2016) and Deudon et al. (2018) who respectively added a reinforcement learning loss and an attention-based encoder. These works were further extended by Kool et al. (2019) and Kwon et al. (2020) to use a general transformer architecture (Vaswani et al., 2017), which has become the standard model choice that we also leverage in this paper. These works have given rise to several variants, improving the architecture (Xin et al., 2021; Luo et al., 2023) or the loss (Kim et al., 2021, 2022; Drakulic et al., 2023; Sun et al., 2024). These improvements are orthogonal to our work and could a priori be combined with our method. It is to be noted that construction approaches are not restricted to routing problems: numerous works have tackled various CO problems, especially on graphs, like Maximum Cut (Dai et al., 2017; Barrett et al., 2020), or Job Shop Scheduling Problem (JSSP) (Zhang et al., 2020; Park et al., 2021).
Improving solutions at inference time
Construction methods make use of a predetermined compute budget to generate one valid solution for a problem instance, depending on the number of necessary action steps. In practice, it is common to have a greater, fixed compute budget to solve the problem at hand such that several trials can be attempted on the same instance. However, continuously rolling out the same learned policy is inefficient as (i) the generated solutions lack diversity (Grinsztajn et al., 2023) and (ii) the information gathered from previous rollouts is not utilized. How to efficiently leverage this extra budget has drawn some attention recently.
The first type of approach focuses on improving the diversity of the generated solutions (Kwon et al., 2020; Grinsztajn et al., 2023; Chalumeau et al., 2023b; Hottung et al., 2024). POMO (Kwon et al., 2020) makes use of different starting points to enable the same policy to generate diverse candidates. Grinsztajn et al. (2023) leverages a population of agents with a loss targeted at specialization on sub-distribution of instances. It was extended in Chalumeau et al. (2023b) and Hottung et al. (2024), respectively replacing the population with a continuous latent space or a discrete context vector.
The second category of methods, which includes MEMENTO, addresses the improvement aspect. These methods are theoretically orthogonal and can be combined with those mentioned earlier. Hottung et al. (2022) employs a parameter-efficient fine-tuning approach on the test instances, whereas Choo et al. (2022) enhances this strategy by incorporating tree search. While these two methods demonstrate impressive performance, they rely on rigid, handcrafted improvement and search procedures, which may not be optimal for diverse problem domains or computational budgets. In contrast, Dernedde et al. (2024) introduces a meta-optimizer that learns to calculate flexible parameter updates based on the reinforce gradient, remaining budget, and best solutions discovered so far. This approach facilitates the learning of budget-dependent adaptation strategies, but is tied to a heatmap-based technique, preventing its applicability to certain CO problems like CVRP.
3 Methods
3.1 Preliminaries
Formulation
A CO problem can be represented as a Markov Decision Process (MDP) denoted by $M=(S,A,R,T,H)$. This formulation encompasses the state space $S$ with states $s_{i}\in S$, the action space $A$ with actions $a_{i}\in A$, the reward function $R(r|s,a)$, the transition function $T(s_{i+1}|s_{i},a_{i})$, and the horizon $H$ indicating the episode duration. Here, the state of a problem instance is portrayed as the sequence of actions taken within that instance, where the subsequent state $s_{t+1}$ is determined by applying the chosen action $a_{t}$ to the current state $s_{t}$. An agent is introduced in the MDP to engage with the problem and seek solutions by learning a policy $\pi(a|s)$. The standard objective considered in RL works is the single shot optimisation, i.e. finding a policy that generates a trajectory $\tau$ which maximises the collected reward: $\pi^{*}=\operatorname*{arg\,max}\nolimits_{\pi}\mathbb{E}_{\tau\sim\pi}[R(\tau)]$, where $R(\tau)=\sum_{t=0}^{H}R(s_{t},a_{t})$.
The specificity of most practical cases in RL for CO is that the policy is given a budget of shots $B$ to find the best possible solution, rather than a single shot. Consequently, the learning objective should rather be : $\pi^{*}=\operatorname*{arg\,max}\nolimits_{\pi}\mathbb{E}_{\tau_{i}\sim\pi}[\max\nolimits_{i=1,...,B}R(\tau_{i})]$. It is hence beneficial to learn policies with higher entropy to provide better exploration, and ideally, policies that can use their previous attempts to (i) explore decisions they have not tried yet (ii) put more focus on promising regions of the space.
3.2 MEMENTO
Adaptation to unseen instances is a desired property for neural solvers. Even when evaluated in the distribution they were trained on, neural solvers are not expected to provide the optimal solution on the first shot, due to the NP-hard property of the problems tackled. Hence, making clever use of the available compute budget for fast online adaptation on specific instances is crucial for performance. Neural solvers could be significantly improved by an approach that can simultaneously use past trajectories to adapt the policy, modify the policy significantly based on few trajectories, and calibrate the exploration/exploitation trade-off based on the available budget.
A way to satisfy the previous desiderata is to store all past attempts in a memory which can be leveraged for subsequent trajectories. This ensures that no information is lost and that promising trajectories can be used more than once. There are many ways of implementing such framework depending on how the information is stored, retrieved, and used in the policy. We would like the memory-based update mechanism to be (i) light-weight to not compromise inference time unduly (ii) agnostic to the underlying model architecture (iii) able to leverage pre-trained memory-less policies (iv) capable of implementing a policy gradient update.
To this end, we introduce MEMENTO, a method that dynamically adapts the action distribution of the neural solver using online collected data. This approach achieves the four desiderata: it stores light-weight data about past attempts, and uses an auxiliary model to update the action logits directly based on the previous outcomes. This allows MEMENTO to be used on top of any pretrained policy and to be totally agnostic to its architecture. In principle, MEMENTO can be combined with most existing construction RL algorithms: Attention-Model (Kool et al., 2019), POMO (Kwon et al., 2020), Poppy  (Grinsztajn et al., 2023), COMPASS (Chalumeau et al., 2023b). It meta-learns the update rule during training, enabling data-efficient adaptation of the policy compared to policy gradient methods. At the end of this section, we show that provided with enough capacity, MEMENTO is able to meta-learn the REINFORCE (Williams, 1992) update. Furthermore, learning the update rule enables MEMENTO to provide the base model with context. Policy gradient updates solely rely on the return and the action probability to compute the update while MEMENTO incorporates additional information such as the remaining adaptation budget.
Figure 1: MEMENTO uses a memory to adapt neural solvers at inference time. When taking a decision, data from similar states is retrieved and prepared (1,2), then processed by a MLP to derive correction logits for each action (3). Summing the original and new logits enables to update the action distribution. The resulting policy is then rolled out (4), and transitions’ data is stored in a memory (5,6), including action taken, log probability, information about the state, and obtained return.
Data stored in memory
The stored data reflects several aspects of the decision process. MEMENTO stores the action and a set of associated scalars that will enable it to decide whether to increase the likelihood of this action in subsequent trajectories: (i) information about the state (ii) log-probability of the action (iii) return of the episode (iv) log-probability of the trajectory.
Processing the memory
The data is retrieved from the memory and processed to create action logits that are summed with the logits output by the base model. In the memory, data rows are organised by nodes. When taking a decision at a specific node, it is unlikely that information from other nodes will be useful, therefore we only retrieve information generated at the same node. To adapt to POMO, we also maintain memories for each starting position, but MEMENTO can be used with or without this aspect. We also append the remaining budget as an additional feature. For each action retrieved from the memory, the associated features are normalized and processed by a Multilayer Perceptron (MLP) $H_{\theta_{M}}$ which outputs a scalar weight. The actions are then averaged based on the weight they receive, which creates the correction logits that will be added to the base policy’s logits. One can view this as a revisited attention gate where the key/query attention is non-linear and the value (the actions) is not projected in a subspace prior to the aggregation.
More formally, let $l=\left[l_{1},l_{2},\dots,l_{n}\right]$ be the logits associated with the $n$ possible actions output by the base policy $\pi_{\theta}$, at a given state $s$. The memory associated with $s$, $M_{s}$, is a sequence of tuples $(a_{i},f_{a_{i}})$, where $a_{i}$ are the past actions tried and $f_{a_{i}}$ are various features associated with the corresponding trajectories, as discussed above. The update is computed by $l_{M}=\sum_{i}\mathbf{a_{i}}H_{\theta_{M}}(f_{a_{i}})$, where $\mathbf{a_{i}}$ is the one-hot encoding of action $a_{i}$. The final logits used to sample the next action are given by $l+l_{M}$.
The intuition behind this architecture choice is that it should be able to learn the REINFORCE update rule. Indeed the REINFORCE loss associated with a new transition is $R\log(\pi_{\theta}(a))$, such that $\frac{\partial R\log(\pi_{\theta}(a))}{\partial l_{a}}=R(1-\pi_{\theta}(a))$. Therefore, simply having $H_{\theta_{M}}(f_{a})$ match $R(1-\pi_{\theta}(a))$ would recover a REINFORCE-like update. This is feasible, as $R$ and $\log(\pi_{\theta}(a))$ are included in the features $f_{a}$. We provide an analysis of the impact of those and compare them to REINFORCE in Section 4.1.
Training
Existing construction methods are trained for one-shot optimization, with a few exceptions like Poppy  (Grinsztajn et al., 2023) and COMPASS (Chalumeau et al., 2023b) that are trained for few-shots. We adapt the training process to take into account the budgeted multi-shot setting in which methods are actually used in practice. We hence rollout the policy as many times as the budget allows on the same instance before taking an update. Our loss is inspired by ECO-DQN (Barrett et al., 2020) and ECORD (Barrett et al., 2022): for a new given trajectory, the reward is the ReLU of the difference between the return and the best return achieved so far on that instance. Hence, those summed rewards equal the best score found over the budget. In practice, to account for the fact that it is harder to get an improvement as we get closer to the end of the budget, we multiply this loss with a weight that increases logarithmically as the budget is consumed. To keep the computations tractable, we still compute a loss at each step, estimate the gradient, and average them sequentially until the budget is reached, at which point we take a gradient update step and consider a new batch of instances.
4 Experiments
We evaluate our method across widely recognized CO problems Travelling Salesman (TSP) and Capacitated Vehicle Routing (CVRP). These problems serve as standard benchmarks for evaluating RL-based CO methods (Deudon et al., 2018; Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023; Hottung et al., 2022; Chalumeau et al., 2023b). In Section 4.1, we validate our method by comparing it to leading approach for single-policy adaptation using online collected data, EAS (Hottung et al., 2022). We present their comparative performance in line with established standard benchmarking from literature, and elucidate the mechanisms employed to adapt the action distribution of the base policy during inference. We then demonstrate how MEMENTO can be combined with SOTA neural solver COMPASS with no additional retraining in Section 4.2. Finally, we scale those methods to larger instances in Section 4.3, outperforming heatmap-based methods.
Code availability
We provide access to the code [^1] Code, checkpoints, and evaluation sets available at https://github.com/instadeepai/memento utilized for training our method and executing all baseline models. Additionally, we release our checkpoints for all problem types and scales, accompanied by the necessary datasets to replicate our findings. We implement our method and experiments in JAX (Bradbury et al., 2018), in a codebase that is compatible with previous JAX implementation released in the community. The two problems are also JAX implementations from Jumanji (Bonnet et al., 2023), enabling us to fully leverage hardware accelerators such as TPUs. Our implementations are optimized for TPU v3-8, aligning with the hardware utilized in our experiments.
Figure 2: Performance improvement of MEMENTO and EAS over a base policy. Results are % relative to a greedy rollout of the given base policy (POMO). The benchmark comprises $8$ datasets with $4$ distinct instance sizes on both TSP (left) and CVRP (right). MEMENTO is consistently outperforming policy-gradient based method EAS, with significant improvement to the base policy.
4.1 Benchmarking MEMENTO against policy gradient fine-tuning
The most direct way to leverage online data for policy update is RL/policy-gradients. By contrast, MEMENTO learns an update based on the collected data. We therefore want to see how those two approaches compare. Since EAS (Hottung et al., 2022) is SOTA method using policy gradient, we benchmark MEMENTO against it on a standard set of instances used in the literature (Kool et al., 2019; Kwon et al., 2020; Hottung et al., 2022). Specifically, for TSP and CVRP, those datasets comprise 10 0001000010\,00010 000 instances drawn from the training distribution. These instances feature the positions of $100$ cities/customers uniformly sampled within the unit square. The benchmark also includes three datasets of distributions not encountered during training, each comprising of $1000$ problem instances with larger sizes: $125$, $150$, and $200$, generated from a uniform distribution across the unit square. We employ the exact same datasets as those utilized in the literature for consistency and comparability.
Setup
For routing problems like TSP and CVRP, POMO (Kwon et al., 2020) is the base single-agent, one-shot architecture that underpins most RL construction solvers. In this set of experiments, MEMENTO and EAS both use POMO as a base policy, and adapt it to get the best possible performance within a given budget. We train MEMENTO until it converges, on the same instance distribution as that used for the initial checkpoint. When assessing active-search performance, each method operates within a fixed budget of $1600$ attempts, a methodology akin to Hottung et al. (2022). In this setup, each attempt comprises of one trajectory per possible starting point. This standardized approach facilitates direct comparison to POMO and EAS, which also utilize rollouts from all starting points at each step. Following recent works, we do not use augmentations with symmetries which have proved to not be critical and greatly increase the cost (Chalumeau et al., 2023b).
Figure 3: Heatmaps of the update rules induced by REINFORCE gradient (left), compared to those learned by MEMENTO (right). Akin to REINFORCE, MEMENTO encourages actions with high returns, particularly when they have low probability. Interestingly, MEMENTO learns an asymmetric rule: requiring the normalised return to be strictly positive to reinforce an action, but encouraging it even more when that condition is reached (higher amplitude for positive updates than negative ones).
Results
The average performance of each method across all problem settings are presented in Fig. 2. The observations we draw are three-fold. First, MEMENTO outperforms its base model (POMO) on the entire benchmark by a significant margin: showing that its adaptive search is superior to stochastic sampling. Second, MEMENTO outperforms EAS on all $8$ tasks (spanning both in- and out-of-distribution) for both environments, highlighting the efficacy of meta-learned policy updates when compared to vanilla policy gradients. Finally, we note that this improvement is significant; for example, on TSP100, MEMENTO is doing $60$% better than sampling, while EAS only $6$%.
To better understand how MEMENTO uses its memory to adapt the action distribution of its base policy, we analyse the logit update that an action gets with respect to its associated data, akin to similar analysis conducted in meta-learning (Lu et al., 2022; Jackson et al., 2024), and compare it to the REINFORCE policy gradient. On Fig. 3, we plot the heatmap of the logit update with respect to the log probability (logp) and return associated with an action.
We observe that the main rules learned by MEMENTO are similar to REINFORCE: an action with low logp and high return gets a positive update while an action with high logp and low return gets discouraged. It is nonetheless interesting to see discrepancies. In particular, MEMENTO’s rules are not symmetric with respect to $x=0$, it only encourages action that are above the mean return, but gives higher amplitude to high-return actions than it discourages low-return ones, pushing for performance while preserving exploration. Note that we also expect some kind of dissymetry from EAS since it combines REINFORCE with a term that increases the likelihood to generate the best solution ever found. Finally, the heatmap shows that the update amplitude provided by MEMENTO is significantly bigger than REINFORCE, explaining the higher data-efficiency observed.
4.2 Zero-shot combination of MEMENTO with new solvers
MEMENTO was designed to be agnostic to the base model, enabling combination with existing and future solvers. In this section, we demonstrate zero-shot combination (no re-training) with state-of-the-art solver COMPASS. COMPASS provides fast adaptation since it uses a collection of diverse pre-trained policies that can be searched to pick the most relevant policy for unseen instances.
Figure 4: Evolution of performance over budget when combining MEMENTO to COMPASS on CVRP200, no re-training needed.
Although providing fast adaptation, it can hardly improve further once the appropriate pre-trained policy has been found. In particular, online collected data cannot be used to update that particular policy’s action distribution. We use MEMENTO’s memory module, trained with POMO as a base policy, and apply it on COMPASS with no additional re-training, and show empirically that we can switch on MEMENTO’s adaptation mechanism during the search and reach a new state-of-the-art on $11$ out of $12$ tasks.
Methodology
We load COMPASS checkpoint from Chalumeau et al. (2023b), and we reuse the checkpoint of MEMENTO that was trained using POMO and combine it with COMPASS. For the first half the budget, we do not use MEMENTO, to let COMPASS find the appropriate pre-trained policy. Once the search starts narrowing down, we activate MEMENTO to adapt the pre-trained policy, with no need to turn off COMPASS’ search. Fig. 4 shows the typical trends observed at inference. We see the typical search reported in Chalumeau et al. (2023b). At $50$% of budget consumption, MEMENTO is activated and one can observe a clear and significant drop in averaged tour length of the latest sampled solutions. The standard deviation also illustrates how the search is suddenly narrowed down.
Table 1: Zero-shot combination of MEMENTO and COMPASS on CVRP. Rules learned by MEMENTO transfer to population-based method COMPASS and achieve SOTA on standard benchmark.
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Obj. Gap Obj. Gap Obj. Gap
COMPASS
MEMENTO (COMPASS)
15.644
15.634
-0.019%
-0.082%
17.511
17.497
0.064%
-0.041%
19.313
19.290
0.485%
0.336%
22.462
22.405
2.098%
1.808%
Results
We evaluate the zero-shot combination of MEMENTO and COMPASS on the standard benchmark reported in Section 4.1. Results for CVRP are presented in Table 1, showing tour length and optimality gap in-distribution (CVRP100) and out-of-distribution. Although requiring no re-training, MEMENTO’s rules combine efficiently with COMPASS and achieves new SOTA. Results on TSP are reported in Appendix A, and results on larger instances (TSP & CVRP) in Section 4.3.
4.3 Scaling RL construction methods to larger instances
Scaling neural solvers to larger instances is a crucial challenge to be tackled in order to bridge the gap to industrial applications. Construction-based solvers are promising approaches, requiring limited expert knowledge whilst providing strong performance. But are still hardly ever trained on larger scales with RL. As a result, their performance reported in the literature is usually quite unfair, being evaluated in large scales instances (>$500$) although having been trained on $100$ nodes. In this section, we explain how we have trained construction solvers POMO and COMPASS on instances of size $500$ with RL (for both TSP and CVRP, checkpoints to be open-sourced); and then use them as new base models to validate properties of MEMENTO at that scale.
RL training on larger instances
Training POMO on instances of size $500$ involves three main steps. First, inspired by curriculum-based methods, we start from a checkpoint pre-trained on TSP100. Second, we use Efficient Attention (Rabe and Staats, 2022) to reduce the memory cost of multi-head attention, enabling to rollout problems in parallel despite the $O(n^{2})$ memory requirement. Third, we use gradient accumulation to keep good estimates despite the constrained smaller instance batch size. These three combined tricks enable to train POMO till convergence within $4$ days, and consequently COMPASS and MEMENTO. We also build the zero-shot combined MEMENTO (COMPASS) checkpoint. We release all checkpoints (TSP & CVRP) publicly to ease future research at this scale.
Figure 5: Performance improvement of MEMENTO over EAS on instances of size $500$ on TSP (left) and CVRP (right) for increasing number of sequential attempts and sizes of attempt batch. Green areas correspond to setting for which MEMENTO’s online adaptation is more efficient than EAS. MEMENTO always outperforms EAS on TSP, and for most settings on CVRP. The difference is particularly significant for low budgets, showing the data-efficiency and robustness of MEMENTO.
MEMENTO on larger instances
On those larger instances, the cost of constructing a solution is significantly increased (for the architecture considered: squared in computation and memory cost, and linear in number of sequential operations). It is also harder to get numerous parallel shots, which most methods depend on. It hence becomes crucial to develop data-efficient methods which stay robust to low data regimes. To assess MEMENTO’s efficiency, we train it on instances of size $500$ (for both TSP and CVRP, checkpoints released) and compare its adaptation to EAS on a set of unseen instances, for a range of budget expressed in number of sequential attempts and size of batched solutions per attempts. Figure 5 reports the percentage of improvement (in absolute cost) brought by MEMENTO over EAS. From this figure, we can observe three main tendencies. First, the data-efficiency of MEMENTO is illustrated by its consistent superiority on low budget (lower left of plots), reaching significant improvement compared to EAS. Then, we can see that, on CVRP, the gap increases with the sequential budget for low parallelism. It demonstrates that the meta-learned rule of MEMENTO stays robust, whereas the gradient estimates of EAS get deteriorated and cannot bring further improvement despite additional sequential budget. To finish with, we can see that MEMENTO fully dominates the heatmap on TSP, even for higher budget. Nevertheless, when budget becomes bigger, EAS is able to outperform MEMENTO, under the condition that batches are large enough.
Table 2: Results of MEMENTO against the baseline algorithms on instances of size $500$ of TSP (a) and CVRP (b) for two budget regimes (low and high). We observe that POMO-based methods are outperforming concurrent methods. MEMENTO achieves single-agent SOTA (resp. overall SOTA) on $3$ out of $4$ settings when combined with POMO (resp. COMPASS).
(a) TSP (n=500)
Method Obj. Gap Time -
Concorde 16.55 $0.000\%$ 38M -
LKH-3 16.55 $0.003\%$ 46M -
Dimes 17.80 $7.55\%$ 2H -
Difusco 17.23 $4.08\%$ 11M -
DeepACO 18.74 $13.23\%$
MOCO 16.84 $1.72\%$
Pointerformer 17.14 $3.56\%$ 1M -
Low Budget High Budget
Method Obj. Gap Time Obj. Gap Time
POMO (sampling)
EAS
COMPASS
MEMENTO (POMO)
MEMENTO (COMPASS)
16.999
16.878
16.811
16.838
16.808
2.736%
2.006%
1.603%
1.766%
1.586%
4M
10M
9M
9M
12M
16.986
16.844
16.800
16.823
16.795
2.659%
1.804%
1.539%
1.673%
1.507%
9M
24M
22M
23M
36M
(b) CVRP (n=500)
Method Obj. Gap Time -
LKH-3 37.229 $0.00\%$ 6H -
Low Budget High Budget
Method Obj. Gap Time Obj. Gap Time
POMO (sampling)
EAS
COMPASS
MEMENTO (POMO)
MEMENTO (COMPASS)
37.501
37.425
37.336
37.367
37.309
0.731%
0.525%
0.287%
0.369%
0.215%
9M
16M
10M
16M
15M
37.456
37.185
37.279
37.306
37.251
0.608%
-0.120%
0.133%
0.206%
0.059%
20M
36M
25M
65M
47M
Results
In Table 2, we report performance of our checkpoints, using datasets introduced in Fu et al. (2021) and commonly used in the literature. For each environment, we report the results on two different setting: (i) Low Budget, where the methods are given 25 0002500025\,00025 000 attempts, (ii) High Budget, where 100 000100000100\,000100 000 attempts are available. We also report results from concurrent RL methods (Qiu et al., 2022; Dernedde et al., 2024; Sun and Yang, 2023; Ye et al., 2023), without method-agnostic local search; and industrial solvers LKH (Helsgaun, 2017) and Concorde (Applegate et al., 2006). Interestingly, stochastically sampling solutions with POMO for less than $10$ minutes already provides competitive results, ranking second among the five leading neural solvers. Adding EAS on top of POMO enables to compete with leading method MOCO on TSP500.
It is the first time that POMO is trained on those instance sizes, making it the first RL-trained neural solver that can be used for large instances both on TSP and CVRP, since most previous large scale RL methods are graph-specific, and hence cannot be applied on CVRP. On this benchmark, MEMENTO achieved state-of-the-art on $3$ of the $4$ tasks when combined with COMPASS, showing significant improvement compared to previous state-of-the-art MOCO. It also proves efficiency on those large scales by being the best single agent method on all TSP instances, and on the low budget CVRP task.
5 Conclusion
We present MEMENTO, a method to improve adaptation of neural CO solvers to unseen instances. MEMENTO enables to condition a policy directly on data collected online during search, facilitating rapid adaptation and efficient use of the search budget. In practice, it proves to outperform stochastic sampling and policy-gradient fine-tuning of the policy, and shows it can be train with one solver, and combined with other existing solvers with no-retraining needed.
We demonstrate the efficacy of MEMENTO by achieving state-of-the-art single-policy adaptation on the standard benchmark on TSP and CVRP, both in and out-of-distribution. Moreover, we find that MEMENTO realises interpretable update rules to the underlying policy; trading off exploration and exploitation over the search budget and rediscovering REINFORCE-style updates. We further demonstrate the flexibility of this approach by achieving zero-shot combination of MEMENTO and COMPASS, achieving state-of-the-art on a benchmark of $12$ tasks. The simplicity and efficacy of this method makes it amenable to a very large set of problems, laying another stone towards the application of neural solvers to real-world problems.
Limitations and future work. MEMENTO incurs additional compute processing and memory-usage when compared to the memory-less base policies which it augments. Practically, we find that in all considered settings, the performance gain outweighs these overheads, however practitioners should ensure that this remains the case in other settings. The adaptation enabled by MEMENTO is determined by the data collected by the base policy, therefore this could limit the efficiency when evaluation instances are very far from the training distribution, or otherwise when the base policy has harmfully poor initial performance. Finally, we believe that the capacity of MEMENTO could be improved if trained with broader distribution, in particular using variable instance sizes.
Acknowledgements
Research supported with Cloud TPUs from Google’s TPU Research Cloud (TRC).
Appendix
[…]
Appendix A Extended results
In Section 4, we compare MEMENTO to popular methods from the literature in numerous settings. In particular, Figure 2 compares using stochastic sampling, MEMENTO and EAS to adapt POMO on the standard benchmark; and Figure 5 compares the relative performances of MEMENTO and EAS on larger instances of TSP and CVRP for several values of batch sizes and budget. In this section, we report additional values and results of other methods for those experiments.
A.1 Standard benchmark
We evaluate our method, MEMENTO, against leading RL methods and industrial solvers. For RL specific methods, we provide results for POMO with greedy action selection and stochastic sampling, and EAS; an active search RL method built on POMO, that fine-tunes a policy on each problem instance. We also report population-based method COMPASS and the zero-shot combination of COMPASS with MEMENTO.
We compare to the heuristic solver LKH3 [Helsgaun, 2017]; the current leading industrial solver of both TSP and CVRP, as well as an exact solver Concorde [Applegate et al., 2006] which is a TSP-specific industrial solver.
We use datasets of 10,000 instances with 100 cities/customer nodes drawn from the training distribution, and three generalization datasets of 1,000 instances of sizes 125, 150, and 200, all from benchmark sets frequently used in the literature [Kool et al., 2019, Kwon et al., 2020, Hottung et al., 2022, Grinsztajn et al., 2023, Chalumeau et al., 2023b]. Tables  3(a) and  3(b) display results for TSP and CVRP on the standard benchmark, respectively. The columns provide the absolute tour length, the optimality gap, and the total inference time for each method.
We can see that in the single-agent setting, MEMENTO leads the entire benchmark, and in the population-based setting, MEMENTO (COMPASS) is leading the whole benchmark. MEMENTO is able to give significant improvement to COMPASS on CVRP, pushing significantly the state-of-the-art on this benchmark. On TSP, the improvement is not significant enough to be visible on the rounded results. To finish with, we can observe that the time cost associated with MEMENTO is reasonable and is worth the performance improvement (except for MEMENTO (COMPASS)).
Table 3: Results of MEMENTO against the baseline algorithms for (a) TSP, (b) CVRP. The methods are evaluated on instances from training distribution as well as on larger instance sizes to test generalization.
(a) TSP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
Concorde 7.765 $0.000\%$ 82M 8.583 $0.000\%$ 12M 9.346 $0.000\%$ 17M 10.687 $0.000\%$ 31M
LKH3 7.765 $0.000\%$ 8H 8.583 $0.000\%$ 73M 9.346 $0.000\%$ 99M 10.687 $0.000\%$ 3H
POMO (greedy) 7.796 0.404% 16S 8.635 0.607% 20S 9.440 1.001% 29S 10.933 2.300% 45S
POMO (sampling) 7.779 0.185% 16S 8.609 0.299% 20S 9.401 0.585% 29S 10.956 2.513% 45S
EAS 7.778 0.161% 39S 8.604 0.238% 46S 9.380 0.363% 64S 10.759 0.672% 91S
MEMENTO (POMO) 7.768 0.045% 43S 8.592 0.109% 52S 9.365 0.202% 77S 10.758 0.664% 115S
COMPASS
MEMENTO (COMPASS)
7.765
7.765
0.008 %
0.008%
20S
45S
8.586
8.586
0.036 %
0.035%
24S
55S
9.354
9.354
0.078%
0.077%
33S
83S
10.724
10.724
0.349%
0.348%
49S
127S
(b) CVRP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
LKH3 15.65 $0.000\%$ 17.50 $0.000\%$ 19.22 $0.000\%$ 22.00 $0.000\%$
POMO (greedy) 15.874 1.430% 24S 17.818 1.818% 34S 19.750 2.757% 52S 23.318 5.992% 87S
POMO (sampling) 15.713 0.399% 24S 17.612 0.642% 34S 19.488 1.393% 52S 23.378 6.264% 87S
EAS 15.663 0.081% 66S 17.536 0.146% 82S 19.321 0.528% 123S 22.541 2.460% 179S
MEMENTO (POMO) 15.657 0.066% 118S 17.521 0.095% 150S 19.317 0.478% 169S 22.492 2.205% 392S
COMPASS
MEMENTO (COMPASS)
15.644
15.634
-0.019%
-0.082%
29S
136S
17.511
17.497
0.064%
-0.041%
39S
162S
19.313
19.290
0.485%
0.336%
56S
180S
22.462
22.405
2.098%
1.808%
85S
460S
A.2 Performance over different number of parallel and sequential attempts
In Section 4.3, we show performance improvement of MEMENTO over EAS on instances of size 500 on TSP and CVRP for increasing number of sequential attempts and size of attempt batch. We report extended results with an additional competitor, POMO. We compare the online adaptation of the methods over four different sizes of batched solutions across increasing sequential attempts for each CO problem. Results from Tables 4(a) and 4(b) show results of the three methods on TSP and CVRP, respectively. The columns show the best tour length performance for various values of sequential attempts expressed as budget, and solution batches of size $N$. Performance is averaged over a set of 128 instances.
Table 4: Results of MEMENTO, POMO and EAS on instances of size 500 of (a) TSP and (b) CVRP for increasing number of sequential attempts and sizes of attempt batch.
(a) TSP
$N=20$ $N=40$
Method 200 400 600 800 1000 200 400 600 800 1000
POMO (sampling) 16.9810 16.9707 16.9638 16.9619 16.9574 16.9717 16.96 16.9549 16.9523 16.9493
EAS 16.9223 16.8923 16.8754 16.864 16.8556 16.8777 16.8468 16.83 16.8208 16.8143
MEMENTO 16.8312 16.81 16.799 16.7939 16.7902 16.8187 16.8018 16.7926 16.7855 16.7815
$N=60$ $N=80$
Method 200 400 600 800 1000 200 400 600 800 1000
POMO (sampling)
EAS
MEMENTO
16.9678
16.8616
16.8129
16.9587
16.8353
16.7966
16.9539
16.8211
16.7867
16.952
16.8134
16.7816
16.9503
16.8084
16.7780
16.9634
16.8521
16.8087
16.9547
16.8234
16.7935
16.9513
16.8106
16.7854
16.9495
16.8014
16.7811
16.9474
16.7941
16.7770
(b) CVRP
$N=20$ $N=40$
Method 200 400 600 800 1000 200 400 600 800 1000
POMO (sampling) 37.5256 37.4917 37.4763 37.4639 37.4570 37.4858 37.4599 37.4475 37.4372 37.4307
EAS 37.6107 37.5988 37.5914 37.582 37.5769 37.5022 37.4546 37.422 37.4017 37.3849
MEMENTO 37.3398 37.2992 37.2731 37.2589 37.2527 37.3172 37.2776 37.2515 37.2357 37.2260
$N=60$ $N=80$
Method 200 400 600 800 1000 200 400 600 800 1000
POMO (sampling)
EAS
MEMENTO
37.4598
37.4569
37.292
37.436
37.3679
37.2607
37.4228
37.3248
37.2305
37.4191
37.2892
37.2171
37.4135
37.2624
37.2092
37.4588
37.3588
37.2887
37.4335
37.2804
37.2556
37.4198
37.2346
37.2382
37.4113
37.2021
37.2268
37.4050
37.1747
37.2142
Appendix B Training procedure
This section describes how an existing pre-trained model can be augmented with MEMENTO, and how we train MEMENTO to acquire adaptation capacities.
Firstly, we take an existing single-agent model that was trained using the REINFORCE algorithm and reuse it as a base model. In our case, the base model is POMO. We augment POMO with a memory module and begin the training procedure which aims to create a policy that is able to use past experiences to make decisions in a multi-shot setting. We initialize the memory module weights with small values such that they barely affect the initial output. Hence, the initial MEMENTO checkpoint maintains the same performance as the pre-trained POMO checkpoint.
In the training procedure, the policy is trained to use data stored in the memory to take decisions and solve a problem instance. This is achieved by training the policy in a budgeted multi-shot setting on a problem instance where past experiences are collected and stored in a memory. The memory is organised by nodes such that only information about a specific node is found in a data row that corresponds to that node. The policy retrieves data from the memory at each budget attempt and learns how to uses the data to decide on its next action.
The details of the MEMENTO training procedure are presented in Algorithm 1 and can be understood as follows. At each iteration, we sample a batch B of instances from a problem distribution $\mathcal{D}$. Then, for each instance $\rho_{i}$ where $i\in 1,\dots,B$, for $K$ budget attempts, we retrieve data from the memory. This is done as follows; given that the current selected node is $a_{j}$, we only retrieve data $M(a_{j})$ associated with node $a_{j}$ and its starting point. However, the method is not agnostic to starting point sampling and would work without it. The features associated with the retrieved action are then normalised. We add remaining budget as an additional feature and process the data by a Multilayer Perceptron (MLP) which outputs a scalar weight for each action. We compute correction logits by averaging the actions based on their respective weights. The correction logits are added to the logits of the base policy. We then rollout the resulting policy $\pi_{\tilde{\theta}}$ on the problem instance (i.e., generate a trajectory which represents a solution to the instance). After every policy rollout attempt, the memory is updated with transitions data such as the action taken, the obtained return, the log-probability of the action and the log-probability of the trajectory. These data are stored in the row that corresponds to the current selected node.
For each problem instance, we want to optimise for the best return. At each attempt, we apply the rectified linear unit (ReLU) function to the difference between the last return and the best return ever obtained. We use the rectified difference to compute the REINFORCE loss at each attempt to avoid having a reward that is too sparse and perform back-propagation through the network parameters of our model (including the encoder, the decoder and the memory networks). The sum of the rectified differences is equal to the best return ever obtained over the budget. As the busget is used, it becomes harder to improve over the previous best, the loss terms hence getting smaller. We found that adding a weight to the terms, with logarithmic increase, helped ensuring that the last terms would not vanish, and thus improved performance.
In practice, we also observe that we can improve performance further by adding an optional refining phase where the base model is frozen, and only the memory module is trained, with a reduced learning rate (multiplied by $0.1$), for a few hours.
Algorithm 1 MEMENTO Training
1:  Input: problem distribution $\mathcal{D}$, problem size $N$, memory $M$, batch size $B$, budget $K$, number of training steps $T$, policy $\mathbb{\pi}_{\theta}$ with pre-trained parameters $\theta$.
2:  initialize memory network parameters $\phi$
3:  combine pre-trained policy parameters and memory network parameters $\tilde{\theta}=(\theta,\phi)$
4:  for step 1 to $T$ do
5:     $\rho_{i}\leftarrow$ Sample($\mathcal{D}$) $\forall i\in 1,\ldots,B$
6:     for attempt 1 to $K$ do
7:        for node 1 to $N$ do
8:           $m_{j}\leftarrow$ Retrieve $(M)$ $\forall j\in 1,\ldots,B$   {Retrieve data from the memory}
9:           $\tau_{i}^{j}\leftarrow$ Rollout $(\rho_{i},\pi_{\tilde{\theta}}(\cdot|m_{j}))$ $\forall i,j\in 1,\ldots,B$
10:        $m_{j}\leftarrow f(m_{j},\tau_{i}^{j})$   {Update the memory with transitions data}
11:        $R_{i}^{*}\leftarrow\max(R_{i}^{*},\mathcal{R}(\tau_{i}^{j}))\ \forall i\in 1,\ldots B$   {Update best solution found so far}
12:        $\nabla L(\tilde{\theta})\leftarrow\frac{1}{\textit{B}}\sum_{i\leq B}\textnormal{{REINFORCE}}(\textnormal{{ReLU}}(\tau_{i}^{j}-R_{i}^{*}))$   {Backpropagate}
13:     $\nabla L(\tilde{\theta})\leftarrow\frac{1}{K}\sum_{i=1}^{K}\nabla L(\tilde{\theta})$
14:     $\tilde{\theta}\leftarrow\tilde{\theta}-\alpha\nabla L(\tilde{\theta})$
Appendix C Hyper-parameters
We report all the hyper-parameters used during train and inference time. For our method MEMENTO, there is no training hyper-parameters to report for instance sizes 125, 150, and 200 as the model used was trained on instances of size 100. The hyper-parameters used for MEMENTO are reported in Table  5. Since we also trained POMO and COMPASS on larger instances, we report hyper-parameters used for POMO and COMPASS in Table  6 and  7, respectively.
Table 5: The hyper-parameters used in MEMENTO
(a) TSP
Phase Hyper-parameters TSP100 TSP(125, 150) TSP200 TSP500
Train time budget 200 - - 200
instances batch size 64 - - 32
starting points 100 - - 30
gradient accumulation steps 200 - - 400
memory size 40 - - 80
number of layers 2 - - 2
hidden layers 8 - - 8
activation GELU - - GELU
learning rate (memory) 0.004 - - 0.004
learning rate (encoder) 0.0001 - - 0.0001
learning rate (decoder) 0.0001 - - 0.0001
Inference time
policy noise
memory size
1
40
0.2
40
0.1
40
0.8
40
(b) CVRP
Phase Hyper-parameters CVRP100 CVRP(125, 150) CVRP200 CVRP500
Train time budget 200 - - 200
instances batch size 64 - - 8
starting points 100 - - 100
gradient accumulation steps 200 - - 800
memory size 40 - - 40
number of layers 2 - - 2
hidden layers 8 - - 8
activation GELU - - GELU
learning rate (memory) 0.004 - - 0.004
learning rate (encoder) 0.0001 - - 0.0001
learning rate (decoder) 0.0001 - - 0.0001
Inference time
policy noise
memory size
0.1
40
0.1
40
0.1
40
0.3
40
Table 6: The hyper-parameters used in POMO
Phase Hyper-parameters TSP500 CVRP500
Train time starting points 200 200
instances batch size 64 32
gradient accumulation steps 1 2
Inference time
policy noise
sampling batch size
1
8
1
8
Table 7: The hyper-parameters used in COMPASS
Phase Hyper-parameters TSP500 CVRP500
Train time latent space dimension 16 16
training sample size 64 32
instances batch size 8 8
gradient accumulation steps 8 16
Inference time
policy noise
num. CMAES components
CMAES init. sigma
sampling batch size
0.5
1
100
8
0.3
1
100
8
Appendix D Model checkpoints
Our experiments focus on two CO routing problems, TSP and CVRP, with methods being trained on two distinct instance sizes: 100 and 500. Whenever possible, we re-use existing checkpoints from the literature; in the remaining cases, we release all our newly trained checkpoints in the repository https://github.com/instadeepai/memento.
We evaluate MEMENTO on two CO problems, TSP and CVRP, and compare the performance to that of two main baselines: POMO [Kwon et al., 2020] and EAS [Hottung et al., 2022]. The checkpoints used to evaluate POMO on TSP and CVRP are the same as the one used in Grinsztajn et al. [2023] and Chalumeau et al. [2023b], and the EAS baseline is executed using the same POMO checkpoint. These checkpoints were taken in the publicly available repository https://github.com/instadeepai/poppy. The POMO checkpoint is used in the initialisation step of MEMENTO (as described in Appendix B). To combine MEMENTO and COMPASS on TSP100, we re-use the COMPASS checkpoint available at https://github.com/instadeepai/compass and add the memory processing layers from the MEMENTO checkpoint trained on TSP100. This checkpoint is also released at https://github.com/instadeepai/memento.
For larger instances, we compare our MEMENTO method to three baselines: POMO, EAS and COMPASS. Since no checkpoint of POMO and COMPASS existed, we trained them with the tricks explained in Section 4. The process to generate the MEMENTO checkpoint and the MEMENTO (COMPASS) checkpoint is then exactly the same. All those checkpoints are available, for both TSP and CVRP.
Appendix E Implementation details
The code-base in written in JAX [Bradbury et al., 2018], and is mostly compatible with recent repositories of neural solvers written in JAX, i.e. Poppy and COMPASS. The problems’ implementation are also written in JAX and fully jittable. Those come from the package Jumanji [Bonnet et al., 2023]. CMA-ES implementation to mix MEMENTO and COMPASS is taken from the research package QDax [Chalumeau et al., 2023a]. Neural networks, optimizers, and many utilities are implemented using the DeepMind JAX ecosystem [DeepMind et al., 2020].